# Privacy Guarantees in Posterior Sampling under Contamination

### v2

@@ Abstract
In recent years differential privacy has been adopted by tech-companies and governmental agencies as the standard for measuring privacy in algorithms. In this article, we study differential privacy in Bayesian posterior sampling settings. We begin by considering differential privacy in the most common privatization setting in which Laplace or Gaussian noise is simply injected into the output. In an effort to achieve better differential privacy, we consider adopting Huber’s contamination model for use within privacy settings, and replace at random data points with samples from a heavy-tailed distribution (instead of injecting noise into the output). We derive bounds for the differential privacy level $(\epsilon,\delta)$ of our approach, without the need to impose the restriction of having a bounded observation and parameter space which is commonly used by existing approaches and literature. We further consider for our approach the effect of sample size on the privacy level and the convergence rate of $(\epsilon,\delta)$ to zero. Asymptotically, our contamination approach is fully private at no cost of information loss. We also provide some examples depicting inference models that our setup is applicable to with a theoretical estimation of the convergence rate, together with some simulations.
62F15,
62J12,
Huber’s Contamination Model,
Posterior Sampling,
Bayesian Inference,
@@ 1 Introduction
With the rapid growth of smart electronic devices in recent decades, the quantity of data generated and collected every day has increased by orders of magnitude. A byproduct of this expansion has led to increased awareness of personal privacy and concerns about privacy leakage as a consequence of utilizing the collected data in studies. In recent years, differential privacy has become of significant interest to computer scientists, statisticians and others for studying the properties and trade-offs in designing algorithms that protect the privacy of individuals, and also being increasingly adopted by tech-companies, for instance, Google (Erlingsson, Pihur and Korolova, 2014; Aktay et al., 2020), Microsoft (Ding, Kulkarni and Yekhanin, 2017; Pereira et al., 2021), Apple (Apple, 2017), or government agencies (Machanavajjhala et al., 2008; Abowd, 2018).
Differential privacy, as defined in Dwork et al. (2006, 2014), asserts the difference in response of a randomized algorithm $\mathcal{A}:\mathcal{X}^{n}\rightarrow\Omega$ when being supplied two neighbouring datasets of size $n$, i.e., one dataset is a modification of the other dataset by changing only one data point. In particular, the response pattern should satisfy the following inequality for any measurable set $S\subset\Omega$ and any neighbouring datasets $\bm{X}$ and $\bm{Z}\in\mathcal{X}^{n}$,
@@ 1.1 Related Literature
In most cases, differential privacy is achieved by purposefully injecting noise into the computation process which perturbs the final outcome. The injection of noise can happen in two places, either at the individual level before data collection or after data collection before releasing the result. The most studied mechanism for privatizing Bayesian inference is the latter, where the outcome is perturbed for privacy. Mainly, there are two types of problems being targeted: direct estimation problems (Foulds et al., 2016; Wang, Fienberg and Smola, 2015; Zhang, Rubinstein and Dimitrakakis, 2016; Bernstein and Sheldon, 2019; Kulkarni et al., 2021) and sampling problems (Dimitrakakis et al., 2013; Zhang, Rubinstein and Dimitrakakis, 2016; Foulds et al., 2016; Heikkilä et al., 2019; Yıldırım and Ermiş, 2019).
@@
The typical approach for privatizing estimation problems is to directly inject Laplacian noise into the steps or the output (Wang, Fienberg and Smola, 2015; Foulds et al., 2016; Zhang, Rubinstein and Dimitrakakis, 2016). Bernstein and Sheldon (2019); Kulkarni et al. (2021) considered the injection of noise after moment approximations of the sufficient statistics to achieve privacy. Gaussian/Laplacian mechanisms suffer from the limitation that the bandwidth of the noise scales with the sensitivity, defined as the maximal difference in response with respect to a pair of neighbouring datasets of the target function. In most applications, the sensitivity is not bounded unless the observation space is bounded.
@@
It is worth noting that the definition of ”Bayesian Differential Privacy” in Triastcyn and Faltings (2019, 2020) describes a different problem compared to the previously mentioned papers studying differential privacy in Bayesian inference following Dwork et al. (2006)’s definition. Triastcyn and Faltings (2019, 2020)’s work treats the dissimilar data point $z$ as a random variable, which is marginalized, as opposed to in (1) where both $\bm{X}$ and $\bm{Z}$ are considered as arbitrary but fixed.
@@
In this paper, we consider the differential privacy property of Bayesian posterior sampling problem $\theta\sim\pi_{n}(\cdot|\bm{X})$ where $\pi_{n}$ is the posterior distribution given the dataset $\bm{X}\in\mathcal{X}^{n}$ of size $n$, modelled by likelihood function $f(x;\theta)$, $\theta\in\Omega$ and prior $\pi_{0}(\theta)$. Thereafter, we will call $\mathcal{X}$ the observation (or data) space where the input data resides and $\Omega$ the parameter space where the model parameter $\theta$ takes values in.
@@
As noted in Dimitrakakis et al. (2013), Bayesian posterior sampling is itself a differentially private mechanism without additional noise under the condition that the log density is Lipschitz continuous in data with respect to some pseudo metric $\rho$ for any fixed parameter $\theta$. The level of privacy can be quantified by the pseudo distance $\rho$ between the dissimilar data points $\rho(\bm{x},\bm{z})$. However, allowing the level of privacy to depend on $\rho(\bm{x},\bm{z})$ implies the algorithm can be arbitrarily non-private when the observation space is unbounded. This limitation is shared by the line of studies following Dimitrakakis et al. (2013) on posterior sampling due to having a similar setup (Wang, Fienberg and Smola, 2015; Foulds et al., 2016; Heikkilä et al., 2019; Yıldırım and Ermiş, 2019).
@@
An additional difficulty in bounding the privacy level of posterior sampling comes from the dependency of sensitivity on the parameter space. The sensitivity of the posterior sampling problem is determined by the sensitivity of the log-posterior density function which depends directly on the sample $\theta$. Thus, the sensitivity could also be unbounded if the parameter space is unbounded (Zhang, Rubinstein and Dimitrakakis, 2016; Foulds et al., 2016; Heikkilä et al., 2019; Yıldırım and Ermiş, 2019; Zhang and Zhang, 2023).
@@
In contrast to adding noise at the outcome level, we consider Huber’s contamination model (Huber, 2004) as a differential privacy mechanism. A body of literature exists in quantifying the trade-off between privacy and efficiency statistical inference under Huber’s contamination model (Rohde and Steinberger, 2020; Cai, Wang and Zhang, 2021; Kroll, 2021; Li, Berrett and Yu, 2022). However, these analyses are local differential privacy models, where the privacy is assessed with respect to datasets of size $1$, whereas we study differential privacy of the whole inference/sampling algorithm under Huber’s contamination model, otherwise known as global differential privacy model.
@@ 1.2 Our contributions
  • We derive the differential privacy properties of a general posterior sampling problem under Huber’s contamination model (Huber, 2004) without assuming bounded parameter space or observation space, which are conditions that often appear explicitly or implicitly in the existing literature to ensure finite sensitivity;
  • We characterise the rate of convergence for the differential privacy cost $(\epsilon,\delta)$ with the number of data points $n$ which, to the best of our knowledge, is yet to be answered;
  • We also show through simulations the method is more robust since the injected noise is independent of the size of the target space (Section 2.3), and introducing some contamination could vastly improve the differential privacy when the original model is differentially private by design (Section 4.2).
@@
$k_{p}(\bm{x};\theta):=(1-p)f(\bm{x};\theta)+pg(x).$
Under a set of mild assumptions on the prior and likelihood function defined later in Section 3.2, we informally state our main result:
Theorem 1.
@@
Let $\bm{X}$,$\bm{Z}$ be neighbouring datasets of size $n$ generated by $\mathbb{P}_{\theta^{*}}$, the true model under contamination, with density function $k_{p_{n}}(\cdot;\theta^{*})$. Under mild assumptions on the set of density functions $k_{p}$ and prior $\pi_{0}$ (Assumptions 1-5 defined in Section 3.2), then there exists a sequence of contamination probabilities $\{p_{n}\}\rightarrow 0$, such that $\forall S$ measurable, $\epsilon,\delta>0$, we have
$\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq\exp(\epsilon)\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\delta,$
$\mathbb{P}_{\theta^{*}}$-almost surelly, where $\mathbb{P}_{\pi_{n}}(S|\bm{X})$ denotes the probability mass of $S$ under posterior $\pi_{n}$ given dataset $\bm{X}$. In addition, the entries of Fisher information also converge to the true model
$I_{p_{n}}(\theta^{*})\rightarrow I(\theta^{*}).$
where $I_{p_{n}}$ is the Fisher information matrix for the contaminated likelihood when the contamination rate is $p_{n}$ and $I(\cdot)=I_{0}(\cdot)$.
@@
Theorem 1 essentially states that, under mild conditions, as the size of the dataset $n\rightarrow\infty$, sampling from the contaminated posterior achieves perfect differential privacy. Additionally, no statistical efficiency is lost since the contamination rate $p_{n}\rightarrow 0$.
1.3 Organization of the Paper
@@
This paper is organized as follows. In Section 2 we introduce the setup of differential privacy and data contamination mechanism, and we demonstrate in Section 2.3 how our approach could improve in statistical efficiency compared with the standard frequentists’ approach to differential privacy in parameter estimation.
@@
Following that, we present our main results on the privacy level of $p$-contaminated posterior sampling in Section 3. Any intermediate results for the main theorems are presented in Appendix A with the longer proofs in the supplementary material (Hu et al., 2024). We aim to lift both restrictions on having bounded parameter space and bounded observation space and derive a bound for privacy level that shrinks with $n$ in Section 3.2, with some discussions and remarks on the assumptions.
@@
Since the link with the $(\epsilon,\delta)$ level and the choice of contamination $p$, $g(\cdot)$ is implicit. To help practitioners tune the model, we present, in Section 4, a simulation-based approach to empirically estimate $\epsilon,\delta$ one gets from a certain model setup, and empirically show that the trace of $(\epsilon,\delta)$ against $n$ follows what is described in Theorem 5 through simulation. On the theory side, we show in Section 5 and in the supplementary material (Appendix C) that our assumptions can be easily satisfied for many regression models when $n$ is large enough. Finally, we conclude the paper with a short discussion in Section 6.
2 Differential Privacy in Bayesian Sampling and Inference
In this section, we first introduce the setup of differential privacy in Bayesian posterior sampling and our contamination mechanism, followed by a comparison with the common frequentist’s approaches to privatising parameter estimation.
@@ 2.1 Differential Privacy in Bayesian Inference
In a typical Bayesian inference setting, we have access to a dataset of size $n$, $\mathcal{D}\in(\mathcal{X})^{n}$ where $\mathcal{X}$ is the space of all possible observations, modelled by a density function $f(\bm{x};\theta):\mathcal{X}\times\Omega\rightarrow\mathbb{R}_{\geq 0}$ with parameter $\theta$. The parameter $\theta$ is a random variable with prior density $\pi_{0}(\theta)$ defined on the measurable space $(\Omega,\mathcal{B}(\Omega))$. Then the posterior distribution for $\theta$ given the dataset $\mathcal{D}$ can be expressed as
@@
  • The likelihood functions $f(\bm{x};\theta)$ and $k_{p}(\bm{x};\theta)$ are jointly continuous in $x$ and $\theta$, and the contamination likelihood $g(x)$ has full support over the observation space;
  • The observation space $\mathcal{X}$ and the parameter space $\Omega$ are both convex subsets of the Euclidean space, but not necessarily bounded or of the same dimension.
For now, we also assume that the model is weakly consistent at $\theta^{*}$ (e.g., Definition 6.1 Ghosal and Van der Vaart (2017)). In the later sections, this weak consistency assumption will be implied by other assumptions.
Definition 1 ($(\epsilon,\delta)$-Differentially Private Posterior Sampling).
The direct sampling from the posterior distribution in (2) is considered to be $(\epsilon,\delta)$-differentially private if for any two datasets $\bm{X},\bm{Z}\in\mathcal{X}^{n}$ which differ by at most one entry, for all $S\in\mathcal{B}(\Omega)$,
$\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq\exp({\epsilon})\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\delta.$ (3)
@@
Two datasets $\bm{X}$ and $\bm{Z}$ are neighbours to each other if and only if $\exists\mathcal{D}$ and $\bm{x},\bm{z}\notin\mathcal{D}$ such that $\bm{X}=\mathcal{D}\cup\{\bm{x}\}$, $\bm{Z}=\mathcal{D}\cup\{\bm{z}\}$.
@@
Definition 2 is perhaps the most adopted definition for neighbouring datasets in the differential privacy literature where the two datasets are explicitly required to have the same cardinality.
In some places, it is alternatively defined as one dataset is a proper subset of the other with strictly one less element, for instance, in Dwork (2008). To avoid ambiguity, we call two datasets $\bm{X}$ and $\bm{Z}$ strictly adjacent if $\bm{Z}=\bm{X}\cup\{\bm{z}\}$ and $\bm{z}\notin\bm{X}$ (or the same statement with $\bm{Z}$ and $\bm{X}$ swapped holds). Our result transports easily to the strictly adjacent case which is discussed in Appendix D in the supplementary material.
Remark 2.
In general, Bayesian posterior sampling is not differentially private. For instance, modelling the likelihood using a Gaussian distribution with known variance $1$ and a Gaussian prior $\mathcal{N}(\mu_{0},1)$ on the mean $\mu$. Let $\bm{X},\bm{Z}$ be neighbouring datasets with $\mathcal{D}=\bm{X}\cap\bm{Z}$, $\bm{X}=\mathcal{D}\cup\{x\}$, $\bm{Z}=\mathcal{D}\cup\{z\}$, where $\mathcal{D}:=\{x_{1},\dots,x_{n}\}$. Then
@@
To improve privacy, usually, some noise is injected into the process. For instance, noise can be injected directly into the posterior distribution to contaminate the output sample. Here, we consider the injection of noise before the inference stage where for any data point $y$ in the dataset $\mathcal{D}$, we replace it with $\mathcal{C}(y)$ by
@@
where $10<p<$ 1, $\bm{x}\sim g(\cdot)$ and $g(\cdot)$ is the density function which has a heavier tail than $f(\bm{x};\theta),\forall\theta$. Thus, the likelihood function of the contaminated dataset is given by $k_{p}(\bm{x};\theta)=(1-p)f(\bm{x};\theta)+pg(\bm{x})$.
@@
We may consider two distinct DP setups with the same contamination strategy. When we discuss the differential privacy of a randomized algorithm $\mathcal{A}$ (see (1), it is important to identify what random process is involved in $\mathcal{A}$, in this case, whether contamination is part of $\mathcal{A}$, i.e., contamination at runtime, or not part of $\mathcal{A}$, i.e., contamination before $\mathcal{A}$. In our setup, the latter is preferred, due to its simplicity. Though, one may also consider the former approach where the uncontaminated neighbouring datasets $\bm{X}$ and $\bm{Z}$ are contaminated at runtime to produce $\mathcal{C}(\bm{X})$ and $\mathcal{C}(\bm{Z})$. Now, $\mathcal{C}(\bm{X})$ and $\mathcal{C}(\bm{Z})$ are random variables and their realizations don’t need to be neighbouring to each other. It is not hard to show that if the pre-contaminated sampling algorithm is $(\epsilon,\delta)$-DP, then dynamic contamination will be at least $(\epsilon,\delta)$-DP. However, since at runtime, $\mathcal{C}(\bm{X})$ and $\mathcal{C}(\bm{Z})$ are not necessarily neighbouring data sets, the resulting problem is more complicated to analyze. Either way, one should treat the contamination density $g(\cdot)$ and the contamination rate $p$ as public knowledge to avoid having model identifiability issues (Mu and Xiong, 2023).

### v3

@@ Abstract
In recent years, differential privacy has been adopted by tech-companies and governmental agencies as the standard for measuring privacy in algorithms. In this article, we study differential privacy in Bayesian posterior sampling settings. We begin by considering differential privacy in the most common privatisation setting in which Laplace or Gaussian noise is injected into the output. In an effort to achieve better differential privacy, we consider adopting Huber’s contamination model for use within privacy settings, and replace at random data points with samples from a heavy-tailed distribution (instead of injecting noise into the output). We derive bounds for the differential privacy level $(\epsilon,\delta)$ of our approach, without requiring bounded observation and parameter spaces, a restriction commonly imposed in the literature. We further consider for our approach the effect of sample size on the privacy level and the rate at which $(\epsilon,\delta)$ converges to zero. Asymptotically, our contamination approach is fully private with no information loss. We also provide examples of inference models for which our approach applies, with theoretical convergence rate analysis and simulation studies.
@@ 1 Introduction
With the rapid growth of smart electronic devices in recent decades, the quantity of data generated and collected every day has increased by orders of magnitude. A byproduct of this expansion is increased awareness of personal privacy and concerns about privacy leakage as a consequence of utilising the collected data in studies. In recent years, differential privacy has become of significant interest to computer scientists, statisticians and others for studying the properties and trade-offs in designing algorithms that protect the privacy of individuals, and is also being increasingly adopted by tech-companies, for instance, Google (Erlingsson et al., 2014; Aktay et al., 2020), Microsoft (Ding et al., 2017; Pereira et al., 2021), Apple (Apple, 2017), or government agencies (Machanavajjhala et al., 2008; Abowd, 2018).
Differential privacy, as defined in Dwork et al. (2006, 2014), bounds the difference in response of a randomized algorithm $\mathcal{A}:\mathcal{X}^{n}\rightarrow\Omega$ when being supplied two neighbouring datasets of size $n$, i.e., one dataset is a modification of the other dataset by changing only one data point. In particular, the response pattern should satisfy the following inequality for any measurable set $S\subset\Omega$ and any neighbouring datasets $\bm{X}$ and $\bm{Z}\in\mathcal{X}^{n}$,
@@ 1.1 Related Literature
In most cases, differential privacy is achieved by purposefully injecting noise into the computation process, which perturbs the final outcome. The injection of noise can happen in two places, either at the individual level before data collection or after data collection before releasing the result. The most studied mechanism for privatising Bayesian inference is the latter, where the outcome is perturbed to ensure privacy. Two main types of problems have been targeted: direct estimation problems (Foulds et al., 2016; Wang et al., 2015; Zhang et al., 2016; Bernstein and Sheldon, 2019; Kulkarni et al., 2021) and sampling problems (Dimitrakakis et al., 2017; Zhang et al., 2016; Foulds et al., 2016; Heikkilä et al., 2019; Yıldırım and Ermiş, 2019).
@@
The typical approach for private estimation problems is to directly inject Laplacian (or Gaussian) noise into intermediate steps or the final output (Wang et al., 2015; Foulds et al., 2016; Zhang et al., 2016). Bernstein and Sheldon (2019); Kulkarni et al. (2021) considered the injection of noise after moment approximations of the sufficient statistics to achieve privacy. Gaussian/Laplacian mechanisms have the limitation that the magnitude of the noise scales with the sensitivity, defined as the maximal difference in the output of the target function with respect to neighbouring datasets. In most applications, the sensitivity is unbounded unless the observation space is bounded.
@@
Note that the definition of ”Bayesian Differential Privacy” in Triastcyn and Faltings (2019, 2020) differs from the previously mentioned work on differential privacy in Bayesian inference following Dwork et al. (2006)’s definition. In Triastcyn and Faltings (2019, 2020), the differing data point $z$ is treated as a random variable and marginalised, in contrast to (1) where both $\bm{X}$ and $\bm{Z}$ are considered as arbitrary but fixed.
@@
In this paper, we consider the differential privacy property of the Bayesian posterior sampling problem $\theta\sim\pi_{n}(\cdot|\bm{X})$ where $\pi_{n}$ is the posterior distribution given the dataset $\bm{X}\in\mathcal{X}^{n}$ of size $n$, modelled by likelihood function $f(x;\theta)$, $\theta\in\Omega$, and prior $\pi_{0}(\theta)$. Thereafter, we will call $\mathcal{X}$ the observation (or data) space where the input data resides and $\Omega$ the parameter space where the model parameter $\theta$ takes values.
@@
As noted in Dimitrakakis et al. (2017), Bayesian posterior sampling is itself a differentially private mechanism without additional noise under the condition that the log density is Lipschitz continuous in data with respect to some pseudometric $\rho$ for any fixed parameter $\theta$. The level of privacy can be quantified by the pseudometric $\rho$ between the differing data points $\rho(\bm{x},\bm{z})$. However, allowing the level of privacy to depend on $\rho(\bm{x},\bm{z})$ implies the algorithm can be arbitrarily non-private when the observation space is unbounded. This limitation is shared by subsequent studies following Dimitrakakis et al. (2017) on posterior sampling due to having a similar setup (Wang et al., 2015; Foulds et al., 2016; Heikkilä et al., 2019; Yıldırım and Ermiş, 2019).
@@
An additional difficulty in bounding the privacy level of posterior sampling comes from the dependence of sensitivity on the parameter space. The sensitivity of the posterior sampling problem is determined by the sensitivity of the log-posterior density function, which depends directly on the sampled parameter $\theta$. Thus, the sensitivity could also be unbounded if the parameter space is unbounded (Zhang et al., 2016; Foulds et al., 2016; Heikkilä et al., 2019; Yıldırım and Ermiş, 2019; Zhang and Zhang, 2023).
@@
In contrast to adding noise at the outcome level, we consider Huber’s contamination model (Huber, 2004) as a differential privacy mechanism. A body of literature exists on quantifying the trade-off between privacy and efficient statistical inference under Huber’s contamination model (Rohde and Steinberger, 2020; Cai et al., 2021; Kroll, 2021; Li et al., 2023). However, these analyses concern local differential privacy models, where privacy is assessed with respect to datasets of size $1$, whereas we study the differential privacy of the entire inference/sampling algorithm under Huber’s contamination model, otherwise known as the central differential privacy model.
Figure 1: Mean Squared Error (MSE) Comparison for Private Mean Estimation Task. $n$ represents the size of the data set, ranging from $100$ to $50000$. MSE is plotted on log10 scale.
1.2 Motivations
Here, we briefly outline the motivation for choosing to privatise the posterior distribution instead of other steps in the inference process. In practice, a broadly applicable approach to achieving differential privacy is to inject Gaussian/Laplace noise at an appropriate step. The magnitude of the noise depends on the sensitivity of the quantity to be privatised. In certain settings, such as logistic regression, the sensitivity of interest is naturally bounded. However, in general, sensitivity is unbounded unless the analysis is restricted to bounded spaces.
When Gaussian/Laplace noise is injected in a straightforward manner, the injected noise usually scales polynomially with the diameter of the bounded domain. If the domain is conservatively chosen to be large, this scaling can result in poor performance. Certain algorithms avoid the need for a bounded-domain assumption by applying appropriate clipping, e.g., Biswas et al. (2020); Huang et al. (2021) for private mean estimation. However, such algorithms are often intricate to design and problem-specific.
In contrast, the Bayesian paradigm provides a general and well-studied foundation for statistical inference. Inference proceeds via the posterior distribution, and sampling from the posterior can naturally be viewed as a “privatisation mechanism” due to its inherent randomness. Although numerous works have investigated differential privacy guarantees for posterior sampling (Wang et al., 2015; Zhang et al., 2016; Foulds et al., 2016; Heikkilä et al., 2019; Yıldırım and Ermiş, 2019; Zhang and Zhang, 2023), most of these results rely on strong assumptions about the likelihood function, which often entail bounded domains for either the dataset $\bm{X}$ or the parameter $\theta$. Given the flexibility of the Bayesian framework, it is of significant interest to develop a general mechanism with theoretical guarantees for privatising posterior distributions without imposing restrictive assumptions. This goal motivates the present work.
@@ 1.3 Our contributions
  • We derive, in a probabilistic setting similar to the work of Hall et al. (2013), the differential privacy properties of a general posterior sampling problem under Huber’s contamination model (Huber, 2004) without assuming bounded parameter space or observation space, which are conditions that often appear explicitly or implicitly in the existing literature to ensure finite sensitivity;
  • We characterise the rate of convergence for the differential privacy cost $(\epsilon,\delta)$ as a function of the number of data points $n$, which, to the best of our knowledge, has yet to be addressed;
  • As an application, we show that the contaminated Bayesian inference may be used to solve private mean estimation problems with comparable performance to state-of-the-art frequentist methods (see Figure 1 and Section 4.3).
@@
$k_{p}(\bm{x};\theta):=(1-p)f(\bm{x};\theta)+pg(\bm{x}).$
We informally state here our main result from Section 3.2:
Result.
@@
Under mild assumptions, there exists a sequence of contamination probabilities $p_{n}\rightarrow 0$ such that $\forall\epsilon,\delta>0$,
$\forall S\text{ measurable, }\,\,\,\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq\exp(\epsilon)\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\delta$
fails with probability shrinking exponentially to 0 as $n\rightarrow\infty$, where the probability is taken with respect to the neighbouring datasets $\bm{X}$ and $\bm{Z}$ (differ by only one entry) generated from the true contaminated density $k_{p_{n}}$.
In addition, for any $\theta$, the Fisher information $I_{p_{n}}(\theta)$ for the contaminated model converges entrywise to that of the uncontaminated model $I(\theta)$
$I_{p_{n}}(\theta)\rightarrow I(\theta).$
@@
The foregoing result demonstrates that, under mild regularity conditions, as the dataset size $n\to\infty$, sampling from the contaminated posterior achieves perfect differential privacy almost surely. Moreover, since the contamination rate satisfies $p_{n}\to 0$, the contaminated posterior distribution converges asymptotically to the true posterior. Consequently, the statistical efficiency of the proposed approach becomes asymptotically equivalent to that of the uncontaminated Bayesian model.
1.4 Organisation of the Paper
@@
This paper is organised as follows. In Section 2, we introduce the setup of differential privacy and the data contamination mechanism, with relevant definitions needed for the main result.
@@
Following that, we present our main results on the privacy level of $p$-contaminated posterior sampling in Section 3. Intermediate results for the main theorems are presented in Appendix A, and the longer proofs appear in the supplementary material (Hu et al., 2026). We lift the restrictions of bounded parameter and observation spaces and derive in Section 3.2 a privacy bound that shrinks with $n$, with accompanying discussions of the assumptions.
@@
The link between the $(\epsilon,\delta)$ level and the contamination parameters $p$ and $g(\cdot)$ is implicit. To help practitioners tune the model, we present, in Section 4, a simulation-based approach to estimate $\epsilon$ (with $\delta$ fixed) for a given model setup, and to examine the privacy level of several regression models as the data size $n$ varies. To illustrate the performance, we compare our approach with two state-of-the-art frequentist methods on a private mean estimation problem, in Section 4.2. On the theoretical side, we show in Section 5 and in the supplementary material (Appendix C) that our assumptions are satisfied for many regression models when $n$ is large enough. Finally, we conclude the paper with a short discussion in Section 6.
2 Preliminaries
In this section, we introduce the basic ideas and definitions needed to construct our framework for establishing differential privacy in Bayesian posterior sampling.
@@ 2.1 Differential Privacy in Bayesian Inference
In a typical Bayesian inference setting, we have access to a dataset of size $n$, $\mathcal{D}\in\mathcal{X}^{n}$, where $\mathcal{X}$ is the observation space and the data are modelled by a density $f(\bm{x};\theta):\mathcal{X}\times\Omega\rightarrow\mathbb{R}_{\geq 0}$ with parameter $\theta\in\Omega$. The parameter $\theta$ is a random variable with prior density $\pi_{0}(\theta)$ on the measurable space $(\Omega,\mathcal{B}(\Omega))$. Then the posterior distribution for $\theta$ given the dataset $\mathcal{D}$ can be expressed as
@@
  • The likelihood functions $f(\bm{x};\theta)$ and $k_{p}(\bm{x};\theta)$ are jointly continuous in $\bm{x}$ and $\theta$, and the contamination likelihood $g(\bm{x})$ has full support over the observation space;
  • The observation space $\mathcal{X}$ and the parameter space $\Omega$ are convex subsets of Euclidean space, but need not be bounded or of the same dimension.
For now, we also assume that the model is weakly consistent at $\theta^{*}$ (see, e.g., Definition 6.1 of Ghosal and Van der Vaart (2017)). In later sections, this weak consistency assumption will be implied by other assumptions.
@@
Two datasets $\bm{X}$ and $\bm{Z}$ are neighbours if and only if $\exists\mathcal{D}$ and $\bm{x},\bm{z}\notin\mathcal{D}$ such that $\bm{X}=\mathcal{D}\cup\{\bm{x}\}$, $\bm{Z}=\mathcal{D}\cup\{\bm{z}\}$.
We adapt the following definition from Hall et al. (2013).
Definition 2 ($(\epsilon_{n},\delta_{n},p_{n})$-Random Differential Privacy).
The direct sampling from the posterior distribution in (2) is considered to be $(\epsilon_{n},\delta_{n},p_{n})$-random differentially private if for any two neighbouring datasets $\bm{X},\bm{Z}\in\mathcal{X}^{n}$, with i.i.d. entries generated from $\mathbb{P}_{\theta^{*}}$,
$\mathbb{P}_{\theta^{*}}\left(\forall S\in\mathcal{B}(\Omega),\,\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq\exp({\epsilon_{n}})\mathbb{P}_{\pi_{n}}(S|\bm{Z})+\delta_{n}\right)\geq 1-p_{n},$ (3)
where $\epsilon_{n},\delta_{n},p_{n}\geq 0$ are decreasing sequences of constants.
@@
Setting $p_{n}=0$ does not imply $(\epsilon_{n},\delta_{n})$-DP, because that would require the inner inequality to be satisfied uniformly for all possible pairs of datasets. Such a condition usually cannot be satisfied uniformly. For instance, when fitting a Bayesian posterior using a Gaussian likelihood of variance $1$ and a Gaussian prior $\mathcal{N}(\mu_{0},1)$ on the mean $\mu$. Let $\bm{X},\bm{Z}$ be neighbouring datasets with $\mathcal{D}=\bm{X}\cap\bm{Z}$, $\bm{X}=\mathcal{D}\cup\{x\}$, $\bm{Z}=\mathcal{D}\cup\{z\}$, where $\mathcal{D}:=\{x_{1},\dots,x_{n}\}$. Then
@@
Here, we consider the injection of noise before the inference stage, where for any data point $\bm{y}$ in the dataset $\mathcal{D}$, we replace it with $\mathcal{C}(\bm{y})$ by
@@
where $0<p<1$, $\bm{x}\sim g(\cdot)$, and $g(\cdot)$ is a density function with a heavier tail than $f(\bm{x};\theta),\forall\theta$. Thus, the likelihood function of the contaminated dataset is given by $k_{p}(\bm{x};\theta)=(1-p)f(\bm{x};\theta)+pg(\bm{x})$.
@@
We may consider two distinct DP setups with the same contamination strategy. When we discuss the differential privacy of a randomised algorithm $\mathcal{A}$, see (1), it is important to identify what random process is involved in $\mathcal{A}$, in this case, whether contamination is part of $\mathcal{A}$, i.e., contamination at runtime, or not part of $\mathcal{A}$, i.e., contamination before $\mathcal{A}$. In our framework, we adopt the latter configuration due to its simplicity.
@@
Nevertheless, one could alternatively consider the former approach, in which uncontaminated neighbouring datasets $\bm{X}$ and $\bm{Z}$ are contaminated at runtime to yield $\mathcal{C}(\bm{X})$ and $\mathcal{C}(\bm{Z})$. In this setting, $\mathcal{C}(\bm{X})$ and $\mathcal{C}(\bm{Z})$ are random variables whose realisations are not necessarily neighbouring datasets. It is not hard to verify that if a pre-contaminated sampling algorithm is $(\epsilon,\delta)$-DP, then contamination at runtime will also satisfy $(\epsilon,\delta)$-DP. However, because $\mathcal{C}(\bm{X})$ and $\mathcal{C}(\bm{Z})$ are not necessarily neighbouring datasets when contamination occurs at runtime, the resulting analysis becomes substantially more complicated. Either way, one should treat the contamination density $g(\cdot)$ and the contamination rate $p$ as public knowledge to avoid having model identifiability issues (Mu and Xiong, 2023).